OpenStudy (anonymous):

Given the sequence 3, 9, 15, which of the following is equivalent to

OpenStudy (anonymous):

54 102 144 168

OpenStudy (anonymous):

t1=3 t2=3+6=9 t3=9+6=15 t4=15+6

OpenStudy (anonymous):

SO 144?

OpenStudy (anonymous):

I think 21.... but not in the option

OpenStudy (anonymous):

huh? which one is it then

OpenStudy (anonymous):

i say 54..

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

what would you say? am i wrong?

OpenStudy (anonymous):

i need to go to bed.. may i know your input?

OpenStudy (anonymous):

Actually what is the question?

OpenStudy (anonymous):

lol look at the top

OpenStudy (anonymous):

Do u want to find the next term?

OpenStudy (anonymous):

you don't understand the question?

OpenStudy (anonymous):

which one of the answer choices is equivalent to the equation.